OpenStudy (anonymous):

For which positive integer values ​​of a is the ratio 36 / (A/10) a) less than 1 b) greater than 36?

OpenStudy (anonymous):

So, the equation \[\frac{ 36 }{\frac{ a }{ 10 }}\] Reduces to:\[36*\frac{ 10 }{ a }\] Which finally reduces to \[\frac{ 360 }{ a }\] And as always with fractions, if the number on the bottom is bigger than the number on the top, the final solution will be less than one, so the answer to a is \[a > 360\] And the answer to b is \[a <10\] because 360/10 is 36 and any number lower would result in a solution higher than 36.

OpenStudy (anonymous):

how did you get \[36\times10\div a\]